In an election results were as follows: 300 people voted for A, 450 people voted for B, and 25 people voted for A and B. If one person is chosen at random from the 1000 eligible voters, find the following probabilities
1. The person voted for A, given that he voted for B.
2. The person voted for B, given that he voted for A.

Answer by stanbon(75887)

In an election results were as follows:
Total # voting = 300 + 450 - 25 = 725
------
300 people voted for A:: P(A) = 300/725
450 people voted for B:: P(B) = 450/725
25 people voted for A and B:: P(A and B) = 25/725
--------------------------------------
If one person is chosen at random from the 1000 eligible voters, find the following probabilities
1. The person voted for A, given that he voted for B.
P(A | B) = P(A and B)/P(B) = 25/450
----------------------------------------------
2. The person voted for B, given that he voted for A.
P(B | A) = P(B and A)/P(A) = 25/300
-----------------
Cheers,
Stan H.
--------